The rate at which cars arrive at the garage is 12 cars per hour. If the arrival of cars at the garage is assumed to follow a poisson distribution, 1 what is the probability that in the next hour there will be at at most 1 car arriving at the garage
1
Expert's answer
2020-10-06T17:50:03-0400

"P(x)=\frac{e^{-12}(12)^{x}}{x!}, x=0,1,2,...\\
P(x\leq1)=P(x=0)+P(x=1)\\
=\frac{e^{-12}(12)^{0}}{0!}+\frac{e^{-12}(12)^{1}}{1!}\approx 0.0000799"